OpenStudy (anonymous):

Solve for x: 3x - 5 = 2x + 6. A. 1 B. -1 C. 11 D. -11

OpenStudy (anonymous):

11

OpenStudy (anonymous):

D 11

OpenStudy (david27):

c

OpenStudy (anonymous):

Sorry C

OpenStudy (anonymous):

what an idiot

OpenStudy (anonymous):

First thing you want to do is get all of the integer values to one side and all x values to the other side. And can everybody PLEASE REFRAIN FROM DIRECT ANSWERS IT IS AGAINST THE CODE OF CONDUCT

OpenStudy (haseeb96):

first bring all the like terms on each sides Look how i do 3x - 5 = 2x + 6 3x - 2x = 6 + 5 x = 11 It is your answer

OpenStudy (anonymous):

thnx guys :)

OpenStudy (haseeb96):

and if you want to check whether my answer is correct or not just put this term in the original equation like we got x=11 and we have the original equation 3x - 5 = 2x + 6 now put x=11 in it 3(11) - 5 = 2(11) + 6 33- 5 = 22+6 28 = 28 if the values are same on both sides it means your answer is correct if not then there must be some error in your answer

OpenStudy (anonymous):

thnx :)